e) A diatomic gas has a density of 1.25 g/L at STP. Calculate its molar mass and identify the gas.


1
Expert's answer
2022-05-30T14:30:04-0400

Under normal conditions, 1 mole of any gas occupies a volume of 22.4 liters. using which we find the molecular mass of the unknown gas. To do this, we increase the density of the gas by 22.4:

Mr = 1,25 * 22,4 = 28 g/mol

Since it is a diatomic gas, we divide the result by 2. resulting in a relative atomic mass: 28 / 2 = 14 gr/mol. The 14 gr mass element in the periodic table of chemical elements is nitrogen

ANSWER: N2 (nitrogen).